# The Hidden DNA of LLM-Generated JavaScript: Structural Patterns Enable High-Accuracy Authorship Attribution

### v1

@@ Abstract.
copyright: nonecopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 2026 International Workshop on Large Language Models for Code, ; April 12, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Supervised learning by classification
@@ 1. Introduction
Authorship attribution (Juola, 2006; Huang et al., 2025), often referred to as stylometry, is the task of determining the author of a document, text, or piece of source code by analyzing stylistic and structural features. In most existing studies on human versus AI authorship, the AI-generated category is represented by a single model (Gurioli et al., 2025; Rahman et al., 2024; Nguyen et al., 2024; Bulla et al., 2024) or, at best, a small set of models (Bukhari et al., 2023), thereby limiting the robustness of the resulting methods (Pan et al., 2024).
In human code attribution, typically only limited samples per author are available (Caliskan-Islam et al., 2015; Abuhamad et al., 2018), while large language models (LLMs) allow the generation of virtually unlimited, task-specific datasets for mapping their coding style. However, because many LLMs are trained on overlapping code corpora (Kocetkov, Denis and Kumar, Divyanshu and Shoeb, Abu Awal Md and others, 2022), they may inherit a broad mixture of stylistic and structural characteristics. With the growing presence of AI-generated content online, researchers have also raised concerns about recursive training and eventual style convergence among models (Shumailov et al., 2023; Mitchell, 2023). If such convergence occurs, model-level attribution, or even model family detection may become increasingly difficult.
Big-tech leaders predict that AI will soon generate most of our source code (Amodei, 2025; Nadella, 2025; Altman, 2025). Thus, robust attribution in LLMs is urgently needed for critical domains like software forensics, malware analysis, cybercrime investigations, and academic integrity (Oman and Cook, 1989). Recent work by Bisztray et al. (Bisztray et al., 2025) shows that different LLMs leave distinguishable fingerprints when writing C code, achieving 94.5% accuracy in telling apart C code generated by five leading LLMs.
@@
In (Bisztray et al., 2025), code stylometry focused on distinguishing C code among five LLM families (Claude-3.5, DeepSeek-v3.1, Gemini-2.5, GPT-4.1, LLaMA-3.3-70B). Extending this line of work, we focus on JavaScript—a flexible, dynamically typed language whose diverse styles and widespread tooling make attribution more challenging. We further investigate whether stylometric attribution is possible within the same model family (GPT-4o, GPT-4o-mini, GPT-4.1, GPT-5, gpt-oss) and explore scaling to larger classification tasks (10–20 classes). To this end, we introduce the first diverse AI-generated JavaScript dataset for authorship attribution and address the following research questions:
@@ 2.1.1. Traditional Machine Learning
Authorship attribution has long been framed as supervised classification, mapping code features into vectors for classifiers such as logistic regression, naïve Bayes, SVMs, or tree ensembles (He et al., 2024). These methods are cheap and interpretable but rely on feature engineering and lag behind deep learning in accuracy (Bisztray et al., 2025).
@@ 2.1.2. Deep Learning
Neural models learn directly from code, capturing stylistic hierarchies that manual features miss (Zafar et al., 2020). Abuhamad et al. achieved 92.3% accuracy on 8,903 authors in Google Code Jam with an RNN+Random Forest hybrid, also showing robustness to obfuscation (Abuhamad et al., 2021). Such models scale better to thousands of authors but require large labeled data.
@@ 2.1.3. Pre-trained and Generative Models
Pre-trained models (i.e., BERT, CodeBERT, RoBERTa) learn contextual embeddings from large code corpora, yielding state-of-the-art attribution with minimal manual features (Feng et al., 2020; Guo et al., 2021; Wang et al., 2021). Fine-tuned CodeT5 variants achieve strong accuracy even across closely related authors (Howard and Ruder, 2018).
Generative LLMs (e.g., GPT-4) support zero/few-shot attribution, reaching 65–69% on hundreds of authors (Choi et al., 2025). Beyond human code, Bukhari et al. distinguished AI- from human-written C code at 92% (Bukhari et al., 2023), while Idialu et al. reported F1 $\approx$ 0.91 for GPT-4 vs. human Python (Idialu et al., 2024), though most treat AI as a single-source class. Bisztray et al. extended this to the AI-vs-AI setting, with a modified CodeT5 surpassing 95% accuracy across five LLMs from different model families (Bisztray et al., 2025).
@@ 2.2.1. Feature Engineering Approaches
Early work fingerprinted authors via lexical cues (keywords, punctuation), formatting habits, and software metrics (Tihanyi et al., 2024). Caliskan-Islam et al. introduced the Code Stylometry Feature Set (CSFS) (Caliskan-Islam et al., 2015). While effective on human subjects, hand-crafted features degrade under obfuscation, adversarial edits, or IDE suggestions (Quiring et al., 2019; Simko et al., 2018; McKnight and Goldberg, 2018).
@@ 2.2.2. Graph and Structural Representations
Graphs capture structural and semantic style. ASTs encode syntax (Guo et al., 2022a), while Program Dependence Graphs (PDGs) and Control Flow Graphs (CFGs) highlight dependencies and control flow. Bogomolov et al. used AST path–based classifiers (PbRF, PbNN) to reach 98% on curated sets but saw sharp drops on realistic corpora (Bogomolov et al., 2021). Graph-based features are more robust to formatting changes than stylometry (Guo et al., 2022b; He et al., 2024).
@@ 2.2.3. Obfuscated Code Attribution
Static stylistic features are fragile under obfuscation (Quiring et al., 2019; Simko et al., 2018; McKnight and Goldberg, 2018), but syntax and AST-based methods remain effective (Caliskan-Islam et al., 2015). One Study achieved $\sim$93% accuracy even on obfuscated code using deep learning  (Abuhamad et al., 2018).
@@ 2.2.4. Execution and Binary-Level Approaches
Dynamic analysis leverages runtime traces (system calls, memory, performance) to reveal algorithmic signatures resistant to superficial edits (Wang et al., 2018). However, sandboxing is costly and mostly limited to malware forensics (Song et al., 2022; Gray et al., 2024; Ferrante et al., 2016). Attribution also extends to binaries: stylistic markers often survive compilation (Rosenblum et al., 2011; He et al., 2024), and despite compiler variability, accuracy remains above chance (Gray et al., 2024).
@@ 2.3. Datasets, Benchmarks and Tasks
Authorship attribution depends on diverse labeled corpora. Widely used datasets for human attribution include Google Code Jam, Codeforces, GitHub OSS, POJ-104, and BigCloneBench (Caliskan-Islam et al., 2015; Alsulami et al., 2017; Abuhamad et al., 2018, 2021; Bogomolov et al., 2021). Yet, classifiers designed for natural language reportedly perform poorly on source code (Pan et al., 2024). For AI vs. human attribution, most studies treat AI as a single-model category (Bukhari et al., 2023; Idialu et al., 2024). Both studies focus exclusively on machine learning methods and report strong results on binary classification of AI vs. human-written C code (F1 $\approx$ 0.91 and Accuracy $\approx$ 92%). However, their evaluation is limited to GPT-4 as the sole source of AI-generated code. The only large-scale AI vs. AI study is by Bisztray et al. (Bisztray et al., 2025), who introduce the LLM-AuthorBench dataset (32k C programs). They achieve 95% accuracy in distinguishing five LLMs, with transformers consistently outperforming traditional ML methods.
@@ 3.2.1. Traditional ML Algorithms
As a baseline, we evaluated five different traditional ML algorithms on stylometric representations of JavaScript code, including $k$-NN, Logistic Regression, Random Forest, Linear SVM, and XGBoost. The pipeline includes pre-processing, feature extraction, and training. Each program was vectorized using TF–IDF over tokens with a vocabulary cap of 400, emphasizing lexical and control-flow patterns known to capture stylistic cues (Caliskan-Islam et al., 2015; Caliskan et al., 2018). Performance was evaluated using accuracy, precision, F1 score, and training time. Since our dataset is balanced, recall does not provide additional information.
@@ 3.2.2. Transformer-based Models
We evaluate three encoder-only architectures (Vaswani et al., 2017), including BERT, CodeBERT, and Longformer. Encoder models are typically favored for classification, while decoder models are preferred for generation with long context.
@@ 3.2.3. Custom CodeT5-JSA Architecture
In  (Bisztray et al., 2025), the authors transformed CodeT5+ (770M parameters) into an encoder-only model by removing the decoder layer and attaching a lightweight classification head (H1) consists of:
@@ 4.1.3. Token length distribution
In practice, trade-offs between context length, attention mechanism, and pretraining objectives determine how effectively stylistic signals are captured. Prior work (Bisztray et al., 2025) finds that representation quality can outweigh raw context window size. For example, a 512-token modified CodeT5 outperforms Longformer (which has a larger context-window).
@@ 4.2.1. Five-Class Attribution for Different Model Families
We first replicate the setup of (Bisztray et al., 2025), where the base BERT model achieved an accuracy of $0.92$ on a five-way attribution task assigning C program code to one of five generator LLMs from distinct model families.
@@ 4.2.4. Different JavaScript Variants
These results extend the observations of (Bisztray et al., 2025) from C to JavaScript, demonstrating that attribution remains feasible both when the number of models scale, and with different transformations. Even as lexical information is deliberately suppressed, model-specific signatures persist through structural and dataflow patterns. Overall, these findings confirm that attribution grounded in semantic and syntactic organization remains tractable and robust even under aggressive obfuscation and fine-grained class expansion.
@@ 5. Limitations and Future Work
  • Language and data diversity: The dataset is restricted to JavaScript, whereas prior work examined C (Bisztray et al., 2025). Extending attribution to other languages (C++, Rust, Python, Java) or to cross-language setups—training on one language and testing on another—would help determine the generality of model fingerprints. Broader datasets covering varied coding domains and prompt styles would also strengthen external validity, as we only focused on web-deveopment.

### v2

@@ Abstract.
copyright: noneccs: Computing methodologies Supervised learning by classification
@@ 1. Introduction
Authorship attribution (authorship; huang_authorship_2025), often referred to as stylometry, is the task of determining the author of a document, text, or piece of source code by analyzing stylistic and structural features. In most existing studies on human versus AI authorship, the AI-generated category is represented by a single model (10992332; rahman_automatic_2024; nguyen_gptsniffer_2024; info15120819) or, at best, a small set of models (10.1145/3605770.3625215), thereby limiting the robustness of the resulting methods (pan2024assessing).
In human code attribution, typically only limited samples per author are available (caliskan-islam_-anonymizing_2015; abuhamad_large-scale_2018), while large language models (LLMs) allow the generation of virtually unlimited, task-specific datasets for mapping their coding style. However, because many LLMs are trained on overlapping code corpora (TheStack2022), they may inherit a broad mixture of stylistic and structural characteristics. With the growing presence of AI-generated content online, researchers have also raised concerns about recursive training and eventual style convergence among models (Shumailov2023CurseOfRecursion; Mitchell2023DataContamination). If such convergence occurs, model-level attribution, or even model family detection may become increasingly difficult.
Big-tech leaders predict that AI will soon generate most of our source code (Amodei2025; Nadella2025; Altman2025). Thus, robust attribution in LLMs is urgently needed for critical domains like software forensics, malware analysis, cybercrime investigations, and academic integrity (OmanCook1989). Recent work by Bisztray et al. (bisztray2025know) shows that different LLMs leave distinguishable fingerprints when writing C code, achieving 94.5% accuracy in telling apart C code generated by five leading LLMs.
@@
In (bisztray2025know), code stylometry focused on distinguishing C code among five LLM families (Claude-3.5, DeepSeek-v3.1, Gemini-2.5, GPT-4.1, LLaMA-3.3-70B). Extending this line of work, we focus on JavaScript—a flexible, dynamically typed language whose diverse styles and widespread tooling make attribution more challenging. We further investigate whether stylometric attribution is possible within the same model family (GPT-4o, GPT-4o-mini, GPT-4.1, GPT-5, gpt-oss) and explore scaling to larger classification tasks (10–20 classes). To this end, we introduce the first diverse AI-generated JavaScript dataset for authorship attribution and address the following research questions:
@@ 2.1.1. Traditional Machine Learning
Authorship attribution has long been framed as supervised classification, mapping code features into vectors for classifiers such as logistic regression, naïve Bayes, SVMs, or tree ensembles (he_authorship_2024). These methods are cheap and interpretable but rely on feature engineering and lag behind deep learning in accuracy (bisztray2025know).
@@ 2.1.2. Deep Learning
Neural models learn directly from code, capturing stylistic hierarchies that manual features miss (zafar_language_2020). Abuhamad et al. achieved 92.3% accuracy on 8,903 authors in Google Code Jam with an RNN+Random Forest hybrid, also showing robustness to obfuscation (abuhamad_large-scale_2021). Such models scale better to thousands of authors but require large labeled data.
@@ 2.1.3. Pre-trained and Generative Models
Pre-trained models (i.e., BERT, CodeBERT, RoBERTa) learn contextual embeddings from large code corpora, yielding state-of-the-art attribution with minimal manual features (feng_codebert_2020; guo_graphcodebert_2021; wang_codet5_2021). Fine-tuned CodeT5 variants achieve strong accuracy even across closely related authors (howard-ruder-2018-universal).
Generative LLMs (e.g., GPT-4) support zero/few-shot attribution, reaching 65–69% on hundreds of authors (choi_i_2025). Beyond human code, Bukhari et al. distinguished AI- from human-written C code at 92% (10.1145/3605770.3625215), while Idialu et al. reported F1 $\approx$ 0.91 for GPT-4 vs. human Python (idialu2024whodunit), though most treat AI as a single-source class. Bisztray et al. extended this to the AI-vs-AI setting, with a modified CodeT5 surpassing 95% accuracy across five LLMs from different model families (bisztray2025know).
@@ 2.2.1. Feature Engineering Approaches
Early work fingerprinted authors via lexical cues (keywords, punctuation), formatting habits, and software metrics (tihanyi_how_2024). Caliskan-Islam et al. introduced the Code Stylometry Feature Set (CSFS) (caliskan-islam_-anonymizing_2015). While effective on human subjects, hand-crafted features degrade under obfuscation, adversarial edits, or IDE suggestions (quiring_misleading_2019; simko2018; mcknight2018).
@@ 2.2.2. Graph and Structural Representations
Graphs capture structural and semantic style. ASTs encode syntax (10.1007/978-981-16-6372-7_70), while Program Dependence Graphs (PDGs) and Control Flow Graphs (CFGs) highlight dependencies and control flow. Bogomolov et al. used AST path–based classifiers (PbRF, PbNN) to reach 98% on curated sets but saw sharp drops on realistic corpora (bogomolov2021authorship). Graph-based features are more robust to formatting changes than stylometry (guo_method_2022; he_authorship_2024).
@@ 2.2.3. Obfuscated Code Attribution
Static stylistic features are fragile under obfuscation (quiring_misleading_2019; simko2018; mcknight2018), but syntax and AST-based methods remain effective (caliskan-islam_-anonymizing_2015). One Study achieved $\sim$93% accuracy even on obfuscated code using deep learning  (abuhamad_large-scale_2018).
@@ 2.2.4. Execution and Binary-Level Approaches
Dynamic analysis leverages runtime traces (system calls, memory, performance) to reveal algorithmic signatures resistant to superficial edits (wang_integration_2018). However, sandboxing is costly and mostly limited to malware forensics (9825799; 10.1145/3653973; 7784595). Attribution also extends to binaries: stylistic markers often survive compilation (rosenblum_who_2011; he_authorship_2024), and despite compiler variability, accuracy remains above chance (10.1145/3653973).
@@ 2.3. Datasets, Benchmarks and Tasks
Authorship attribution depends on diverse labeled corpora. Widely used datasets for human attribution include Google Code Jam, Codeforces, GitHub OSS, POJ-104, and BigCloneBench (caliskan-islam_-anonymizing_2015; alsulami_source_2017; abuhamad_large-scale_2018; abuhamad_large-scale_2021; bogomolov2021authorship). Yet, classifiers designed for natural language reportedly perform poorly on source code (pan2024assessing). For AI vs. human attribution, most studies treat AI as a single-model category (10.1145/3605770.3625215; idialu2024whodunit). Both studies focus exclusively on machine learning methods and report strong results on binary classification of AI vs. human-written C code (F1 $\approx$ 0.91 and Accuracy $\approx$ 92%). However, their evaluation is limited to GPT-4 as the sole source of AI-generated code. The only large-scale AI vs. AI study is by Bisztray et al. (bisztray2025know), who introduce the LLM-AuthorBench dataset (32k C programs). They achieve 95% accuracy in distinguishing five LLMs, with transformers consistently outperforming traditional ML methods.
@@ 3.2.1. Traditional ML Algorithms
As a baseline, we evaluated five different traditional ML algorithms on stylometric representations of JavaScript code, including $k$-NN, Logistic Regression, Random Forest, Linear SVM, and XGBoost. The pipeline includes pre-processing, feature extraction, and training. Each program was vectorized using TF–IDF over tokens with a vocabulary cap of 400, emphasizing lexical and control-flow patterns known to capture stylistic cues (caliskan-islam_-anonymizing_2015; caliskan_when_2018). Performance was evaluated using accuracy, precision, F1 score, and training time. Since our dataset is balanced, recall does not provide additional information.
@@ 3.2.2. Transformer-based Models
We evaluate three encoder-only architectures (vaswani_attention_2017), including BERT, CodeBERT, and Longformer. Encoder models are typically favored for classification, while decoder models are preferred for generation with long context.
@@ 3.2.3. Custom CodeT5-JSA Architecture
In  (bisztray2025know), the authors transformed CodeT5+ (770M parameters) into an encoder-only model by removing the decoder layer and attaching a lightweight classification head (H1) consists of:
@@ 4.1.3. Token length distribution
In practice, trade-offs between context length, attention mechanism, and pretraining objectives determine how effectively stylistic signals are captured. Prior work (bisztray2025know) finds that representation quality can outweigh raw context window size. For example, a 512-token modified CodeT5 outperforms Longformer (which has a larger context-window).
@@ 4.2.1. Five-Class Attribution for Different Model Families
We first replicate the setup of (bisztray2025know), where the base BERT model achieved an accuracy of $0.92$ on a five-way attribution task assigning C program code to one of five generator LLMs from distinct model families.
@@ 4.2.4. Different JavaScript Variants
These results extend the observations of (bisztray2025know) from C to JavaScript, demonstrating that attribution remains feasible both when the number of models scale, and with different transformations. Even as lexical information is deliberately suppressed, model-specific signatures persist through structural and dataflow patterns. Overall, these findings confirm that attribution grounded in semantic and syntactic organization remains tractable and robust even under aggressive obfuscation and fine-grained class expansion.
@@ 5. Limitations and Future Work
  • Language and data diversity: The dataset is restricted to JavaScript, whereas prior work examined C (bisztray2025know). Extending attribution to other languages (C++, Rust, Python, Java) or to cross-language setups—training on one language and testing on another—would help determine the generality of model fingerprints. Broader datasets covering varied coding domains and prompt styles would also strengthen external validity, as we only focused on web-deveopment.